# On the Lipschitz Continuity of Set Aggregation Functions and Neural Networks for Sets

### v2

@@ A.5.2 The $\textsc{NN}_{\textsc{mean}}$ model is Lipschitz continuous with respect to EMD
Let $X=\{\!\!\{\mathbf{v}_{1},\ldots,\mathbf{v}_{m}\}\!\!\}$, $Y=\{\!\!\{\mathbf{u}_{1},\ldots,\mathbf{u}_{n}\}\!\!\}$ be two multisets of vectors. Let also $\text{Lip}(f_{\text{MLP}_{1}})$ and $\text{Lip}(f_{\text{MLP}_{2}})$ denote the Lipschitz constants of the $f_{\text{MLP}_{1}}$ and $f_{\text{MLP}_{2}}$, respectively. Finally, let $\mathbf{F}^{*}$ denote the matrix that minimizes $d_{\text{EMD}}(X,Y)$. Then, we have:
$\bigg{|}\bigg{|}f_{\text{MLP}_{2}}\bigg{(}\frac{1}{m}\sum_{i=1}^{m}f_{\text{MLP}_{1}}(\mathbf{v}_{i})\bigg{)}-f_{\text{MLP}_{2}}\bigg{(}\frac{1}{n}\sum_{j=1}^{n}f_{\text{MLP}_{1}}(\mathbf{u}_{j})\bigg{)}\bigg{|}\bigg{|}$ $\leq\text{Lip}(f_{\text{MLP}_{2}})\bigg{|}\bigg{|}\frac{1}{m}\sum_{i=1}^{m}f_{\text{MLP}_{1}}(\mathbf{v}_{i})-\frac{1}{n}\sum_{j=1}^{n}f_{\text{MLP}_{1}}(\mathbf{u}_{j})\bigg{|}\bigg{|}$
$\leq\text{Lip}(f_{\text{MLP}_{2}})\sum_{i=1}^{m}\sum_{j=1}^{n}[\mathbf{F}]_{ij}||f_{\text{MLP}_{1}}(\mathbf{v}_{i})-f_{\text{MLP}_{1}}(\mathbf{v}_{j})||$
$\leq\text{Lip}(f_{\text{MLP}_{2}})\sum_{i=1}^{m}\sum_{j=1}^{n}[\mathbf{F}^{*}]_{ij}||f_{\text{MLP}_{1}}(\mathbf{v}_{i})-f_{\text{MLP}_{1}}(\mathbf{v}_{j})||$
$\leq\text{Lip}(f_{\text{MLP}_{2}})\sum_{i=1}^{m}\sum_{j=1}^{n}[\mathbf{F}^{*}]_{ij}\text{Lip}(f_{\text{MLP}_{1}})||\mathbf{v}_{i}-\mathbf{v}_{j}||$
$=\text{Lip}(f_{\text{MLP}_{1}})\text{Lip}(f_{\text{MLP}_{2}})\sum_{i=1}^{m}\sum_{j=1}^{n}[\mathbf{F}^{*}]_{ij}||\mathbf{v}_{i}-\mathbf{v}_{j}||$
@@
A.5.3 The $\textsc{NN}_{\textsc{max}}$ model is Lipschitz continuous with respect to the Hausdorff distance
Let $X=\{\!\!\{\mathbf{v}_{1},\ldots,\mathbf{v}_{m}\}\!\!\}$, $Y=\{\!\!\{\mathbf{u}_{1},\ldots,\mathbf{u}_{n}\}\!\!\}$ be two multisets of vectors. Let also $\text{Lip}(f_{\text{MLP}_{1}})$ and $\text{Lip}(f_{\text{MLP}_{2}})$ denote the Lipschitz constants of the $f_{\text{MLP}_{1}}$ and $f_{\text{MLP}_{2}}$, respectively. Let $f_{\text{MLP}_{1}}(X)=\big{\{}\!\!\big{\{}f_{\text{MLP}_{1}}(\mathbf{v}_{1}),\ldots,f_{\text{MLP}_{1}}(\mathbf{v}_{m})\big{\}}\!\!\big{\}}$ and $f_{\text{MLP}_{1}}(Y)=\big{\{}\!\!\big{\{}f_{\text{MLP}_{1}}(\mathbf{u}_{1}),\ldots,f_{\text{MLP}_{1}}(\mathbf{u}_{n})\big{\}}\!\!\big{\}}$. Let also $\mathbf{v}_{i^{*}},\mathbf{u}_{j^{*}}$ and $f_{\text{MLP}_{1}}(\mathbf{v}_{i^{**}}),f_{\text{MLP}_{1}}(\mathbf{u}_{j^{**}})$ denote the vectors from which $d_{H}(X,Y)$ and $d_{H}\Big{(}f_{\text{MLP}_{1}}(X),f_{\text{MLP}_{1}}(Y)\Big{)}$ emerge, i. e., $d_{H}(X,Y)=||\mathbf{v}_{i^{*}}-\mathbf{u}_{j^{*}}||$ and $d_{H}\Big{(}f_{\text{MLP}_{1}}(X),f_{\text{MLP}_{1}}(Y)\Big{)}=\big{|}\big{|}f_{\text{MLP}_{1}}(\mathbf{v}_{i^{**}})-f_{\text{MLP}_{1}}(\mathbf{u}_{j^{**}})\big{|}\big{|}$. Without loss of generality, we also assume that $h\Big{(}f_{\text{MLP}_{1}}(X),f_{\text{MLP}_{1}}(Y)\Big{)}\geq h\Big{(}f_{\text{MLP}_{1}}(Y),f_{\text{MLP}_{1}}(X)\Big{)}$. Then, we have:
$\bigg{|}\bigg{|}f_{\text{MLP}_{2}}\Big{(}\big{(}f_{\text{MLP}_{1}}(\mathbf{v}_{i})\big{)}_{\text{max}}\Big{)}-f_{\text{MLP}_{2}}\Big{(}\big{(}f_{\text{MLP}_{1}}(\mathbf{u}_{j})\big{)}_{\text{max}}\Big{)}\bigg{|}\bigg{|}$ $\leq\text{Lip}(f_{\text{MLP}_{2}})\Big{|}\Big{|}\big{(}f_{\text{MLP}_{1}}(\mathbf{v}_{i})\big{)}_{\text{max}}-\big{(}f_{\text{MLP}_{1}}(\mathbf{u}_{j})\big{)}_{\text{max}}\Big{|}\Big{|}$
$\leq\text{Lip}(f_{\text{MLP}_{2}})\sqrt{d}\,\,d_{H}\Big{(}f_{\text{MLP}_{1}}(X),f_{\text{MLP}_{1}}(Y)\Big{)}$
$=\sqrt{d}\text{Lip}(f_{\text{MLP}_{2}})\max_{i\in[m]}\min_{j\in[n]}\Big{|}\Big{|}f_{\text{MLP}_{1}}(\mathbf{v}_{i})-f_{\text{MLP}_{1}}(\mathbf{u}_{j})\Big{|}\Big{|}$
$\leq\sqrt{d}\text{Lip}(f_{\text{MLP}_{2}})\text{Lip}(f_{\text{MLP}_{1}})\max_{i\in[m]}\min_{j\in[n]}||\mathbf{v}_{i}-\mathbf{u}_{j}||$
@@ A.5.3 The $\textsc{NN}_{\textsc{max}}$ model is Lipschitz continuous with respect to the Hausdorff distance
A.6 Proof of Lemma 3.4
A.6.1 The $\textsc{NN}_{\textsc{mean}}$ model is Lipschitz continuous with respect to the matching distance
Let $\mathcal{X}$ denote a set that contains multisets of vectors of equal cardinalities, i. e., $|X|=M$ and $X\in\mathcal{S}(\mathbb{R}^{d}\setminus\{\mathbf{0}\})$, $\forall X\in\mathcal{X}$ where $M\in\mathbb{N}$. Let $X,Y\in\mathcal{X}$ denote two multisets. We have shown in subsection A.5.2 above that:
$\bigg{|}\bigg{|}f_{\text{MLP}_{2}}\bigg{(}\frac{1}{m}\sum_{i=1}^{m}f_{\text{MLP}_{1}}(\mathbf{v}_{i})\bigg{)}-f_{\text{MLP}_{2}}\bigg{(}\frac{1}{n}\sum_{j=1}^{n}f_{\text{MLP}_{1}}(\mathbf{u}_{j})\bigg{)}\bigg{|}\bigg{|}$ $\leq\text{Lip}(f_{\text{MLP}_{1}})\text{Lip}(f_{\text{MLP}_{2}})d_{\text{EMD}}(X,Y)$
@@ A.6.1 The $\textsc{NN}_{\textsc{mean}}$ model is Lipschitz continuous with respect to the matching distance
$\bigg{|}\bigg{|}f_{\text{MLP}_{2}}\bigg{(}\frac{1}{m}\sum_{i=1}^{m}f_{\text{MLP}_{1}}(\mathbf{v}_{i})\bigg{)}-f_{\text{MLP}_{2}}\bigg{(}\frac{1}{n}\sum_{j=1}^{n}f_{\text{MLP}_{1}}(\mathbf{u}_{j})\bigg{)}\bigg{|}\bigg{|}$ $\leq\text{Lip}(f_{\text{MLP}_{1}})\text{Lip}(f_{\text{MLP}_{2}})d_{\text{EMD}}(X,Y)$
@@
A.6.2 The $\textsc{NN}_{\textsc{sum}}$ model is Lipschitz continuous with respect to the matching distance
Let $\mathcal{X}$ denote a set that contains multisets of vectors of equal cardinalities, i. e., $|X|=M$ and $X\in\mathcal{S}(\mathbb{R}^{d}\setminus\{\mathbf{0}\})$, $\forall X\in\mathcal{X}$ where $M\in\mathbb{N}$. Let $X,Y\in\mathcal{X}$ denote two multisets. Let $\pi^{*}$ denote the matching produced by the solution of the matching distance function:
$\pi^{*}=\operatorname*{arg\,min}_{\pi\in\mathfrak{S}_{M}}\sum_{i=1}^{M}||\mathbf{v}_{\pi(i)}-\mathbf{u}_{i}||\$
@@ A.6.2 The $\textsc{NN}_{\textsc{sum}}$ model is Lipschitz continuous with respect to the matching distance
$\bigg{|}\bigg{|}f_{\text{MLP}_{2}}\bigg{(}\sum_{i=1}^{M}f_{\text{MLP}_{1}}(\mathbf{v}_{i})\bigg{)}-f_{\text{MLP}_{2}}\bigg{(}\sum_{j=1}^{M}f_{\text{MLP}_{1}}(\mathbf{u}_{j})\bigg{)}\bigg{|}\bigg{|}$ $\leq\text{Lip}(f_{\text{MLP}_{2}})\bigg{|}\bigg{|}\sum_{i=1}^{M}f_{\text{MLP}_{1}}(\mathbf{v}_{i})-\sum_{j=1}^{M}f_{\text{MLP}_{1}}(\mathbf{u}_{j})\bigg{|}\bigg{|}$
[…]
$\leq\text{Lip}(f_{\text{MLP}_{2}})\sum_{i=1}^{M}\text{Lip}(f_{\text{MLP}_{1}})||\mathbf{v}_{\pi^{*}(i)}-\mathbf{u}_{i}||$
$\leq\text{Lip}(f_{\text{MLP}_{1}})\text{Lip}(f_{\text{MLP}_{2}})\sum_{i=1}^{M}||\mathbf{v}_{\pi^{*}(i)}-\mathbf{u}_{i}||$
@@
which concludes the proof. The $\textsc{NN}_{\textsc{sum}}$ model is thus Lipschitz continuous with respect to the matching distance and its Lipschitz constant is upper bounded by $\text{Lip}(f_{\text{MLP}_{1}})\text{Lip}(f_{\text{MLP}_{2}})$.
A.6.3 The $\textsc{NN}_{\textsc{sum}}$ model is Lipschitz continuous with respect to the matching distance
Let $\mathcal{X}$ denote a set that contains multisets of vectors of equal cardinalities, i. e., $|X|=M$ and $X\in\mathcal{S}(\mathbb{R}^{d})$, $\forall X\in\mathcal{X}$ where $M\in\mathbb{N}$. Let $X,Y\in\mathcal{X}$ denote two multisets. We have shown in subsection A.6.2 above that:
$\bigg{|}\bigg{|}f_{\text{MLP}_{2}}\bigg{(}\sum_{i=1}^{M}f_{\text{MLP}_{1}}(\mathbf{v}_{i})\bigg{)}-f_{\text{MLP}_{2}}\bigg{(}\sum_{j=1}^{M}f_{\text{MLP}_{1}}(\mathbf{u}_{j})\bigg{)}\bigg{|}\bigg{|}$ $\leq\text{Lip}(f_{\text{MLP}_{1}})\text{Lip}(f_{\text{MLP}_{2}})d_{M}(X,Y)$

### v3

@@ B.7.1 The $\textsc{NN}_{\textsc{mean}}$ model is Lipschitz continuous with respect to EMD
Let $X=\{\!\!\{\mathbf{v}_{1},\ldots,\mathbf{v}_{m}\}\!\!\}$, $Y=\{\!\!\{\mathbf{u}_{1},\ldots,\mathbf{u}_{n}\}\!\!\}$ be two multisets of vectors. Let also $\text{Lip}(f_{\text{MLP}_{1}})$ and $\text{Lip}(f_{\text{MLP}_{2}})$ denote the Lipschitz constants of $f_{\text{MLP}_{1}}$ and $f_{\text{MLP}_{2}}$, respectively. Finally, let $\mathbf{F}^{*}$ denote the matrix that minimizes $d_{\text{EMD}}(X,Y)$. Then, we have:
$\bigg\|f_{\text{MLP}_{2}}\bigg(\frac{1}{m}\sum_{i=1}^{m}f_{\text{MLP}_{1}}(\mathbf{v}_{i})\bigg)-f_{\text{MLP}_{2}}\bigg(\frac{1}{n}\sum_{j=1}^{n}f_{\text{MLP}_{1}}(\mathbf{u}_{j})\bigg)\bigg\|$ $\leq\text{Lip}(f_{\text{MLP}_{2}})\bigg\|\frac{1}{m}\sum_{i=1}^{m}f_{\text{MLP}_{1}}(\mathbf{v}_{i})-\frac{1}{n}\sum_{j=1}^{n}f_{\text{MLP}_{1}}(\mathbf{u}_{j})\bigg\|$
$\leq\text{Lip}(f_{\text{MLP}_{2}})\sum_{i=1}^{m}\sum_{j=1}^{n}[\mathbf{F}^{*}]_{ij}\|f_{\text{MLP}_{1}}(\mathbf{v}_{i})-f_{\text{MLP}_{1}}(\mathbf{v}_{j})\|$
[…]
$\leq\text{Lip}(f_{\text{MLP}_{2}})\sum_{i=1}^{m}\sum_{j=1}^{n}[\mathbf{F}^{*}]_{ij}\text{Lip}(f_{\text{MLP}_{1}})\|\mathbf{v}_{i}-\mathbf{v}_{j}\|$
$=\text{Lip}(f_{\text{MLP}_{1}})\text{Lip}(f_{\text{MLP}_{2}})\sum_{i=1}^{m}\sum_{j=1}^{n}[\mathbf{F}^{*}]_{ij}\|\mathbf{v}_{i}-\mathbf{v}_{j}\|$
@@
B.7.2 There exist $\textsc{NN}_{\textsc{sum}}$ models which are not Lipschitz continuous with respect to the matching distance
$\mathbf{v}_{X}=f_{2}\bigg(\text{ReLU}\Big(\big\{\!\!\big\{f_{1}(v_{1}),\ldots,f_{1}(v_{m})\big\}\!\!\big\}\Big)\bigg)$
where $X=\{\!\!\{v_{1},\ldots,v_{m}\}\!\!\}$ is a multiset of scalars, and $f_{1}$ and $f_{2}$ are fully-connected layers, i. e., $f_{1}(x)=a_{1}\,x+b_{1}$ and $f_{2}(x)=a_{2}\,x+b_{2}$. Furthermore, suppose that $a_{1}>0$, $b_{1}>0$ and $a_{2}>0$. Note that the Lipschitz constants of $f_{1}$ and $f_{2}$ are $\text{Lip}(f_{1})=a_{1}$ and $\text{Lip}(f_{2})=a_{2}$, respectively.
$\bigg|f_{2}\bigg(\sum_{j=1}^{2}f_{1}(c)\bigg)-f_{2}\bigg(f_{1}(c)\bigg)\bigg|$ $=\bigg|f_{2}\bigg(\sum_{j=1}^{2}\text{ReLU}(a_{1}\,c+b_{1})\bigg)-f_{2}\bigg(\text{ReLU}(a_{1}\,c+b_{1})\bigg)\bigg|$
$=\Big|f_{2}\Big(2\,(a_{1}\,c+b_{1})\Big)-f_{2}\Big(a_{1}\,c+b_{1}\Big)\Big|$
$=\Big|a_{2}\Big(2\,(a_{1}\,c+b_{1})\Big)+b_{2}-a_{2}\Big(a_{1}\,c+b_{1}\Big)-b_{2}\Big|$
$=|a_{2}(a_{1}\,c+b_{1})|$
$>|a_{2}(a_{1}\,c+c\,L\,a_{1})|$
$=|(L+1)\,a_{2}\,a_{1}\,c|$
$=(L+1)\,\text{Lip}(f_{2})\,\text{Lip}(f_{1})\,c$
$>L\,\text{Lip}(f_{2})\,\text{Lip}(f_{1})\,d_{M}(X,Y)$
Therefore, for any $L>0$, there exist $X,Y\in\mathcal{S}(\mathbb{R}^{d})$ such that $\Big\|f_{2}\Big(\sum_{\mathbf{v}\in X}f_{1}(\mathbf{v})\Big)-f_{2}\Big(\sum_{\mathbf{u}\in Y}f_{1}(\mathbf{u})\Big)\Big\|>L\,\text{Lip}(f_{2})\,\text{Lip}(f_{1})\,d_{M}(X,Y)$. Based on the above, there exist $\textsc{NN}_{\textsc{sum}}$ models which are not Lipschitz continuous with respect to the matching distance.
B.7.3 The $\textsc{NN}_{\textsc{max}}$ model is Lipschitz continuous with respect to the Hausdorff distance
Let $X=\{\!\!\{\mathbf{v}_{1},\ldots,\mathbf{v}_{m}\}\!\!\}$, $Y=\{\!\!\{\mathbf{u}_{1},\ldots,\mathbf{u}_{n}\}\!\!\}$ be two multisets of vectors. Let also $\text{Lip}(f_{\text{MLP}_{1}})$ and $\text{Lip}(f_{\text{MLP}_{2}})$ denote the Lipschitz constants of $f_{\text{MLP}_{1}}$ and $f_{\text{MLP}_{2}}$, respectively. Let $f_{\text{MLP}_{1}}(X)=\big\{\!\!\big\{f_{\text{MLP}_{1}}(\mathbf{v}_{1}),\ldots,f_{\text{MLP}_{1}}(\mathbf{v}_{m})\big\}\!\!\big\}$ and $f_{\text{MLP}_{1}}(Y)=\big\{\!\!\big\{f_{\text{MLP}_{1}}(\mathbf{u}_{1}),\ldots,f_{\text{MLP}_{1}}(\mathbf{u}_{n})\big\}\!\!\big\}$. Let also $\big(f_{\text{MLP}_{1}}(X)\big)_{\text{max}}$ and $\big(f_{\text{MLP}_{1}}(Y)\big)_{\text{max}}$ denote the output of the max operator applied to multisets $f_{\text{MLP}_{1}}(X)$ and $f_{\text{MLP}_{1}}(Y)$, respectively. Without loss of generality, we also assume that $h\Big(f_{\text{MLP}_{1}}(X),f_{\text{MLP}_{1}}(Y)\Big)\geq h\Big(f_{\text{MLP}_{1}}(Y),f_{\text{MLP}_{1}}(X)\Big)$. Then, we have:
$\bigg\|f_{\text{MLP}_{2}}\Big(\big(f_{\text{MLP}_{1}}(X)\big)_{\text{max}}\Big)-f_{\text{MLP}_{2}}\Big(\big(f_{\text{MLP}_{1}}(Y)\big)_{\text{max}}\Big)\bigg\|$ $\leq\text{Lip}(f_{\text{MLP}_{2}})\Big\|\big(f_{\text{MLP}_{1}}(\mathbf{v}_{i})\big)_{\text{max}}-\big(f_{\text{MLP}_{1}}(\mathbf{u}_{j})\big)_{\text{max}}\Big\|$
[…]
$=\sqrt{d}\text{Lip}(f_{\text{MLP}_{2}})\max_{i\in[m]}\min_{j\in[n]}\Big\|f_{\text{MLP}_{1}}(\mathbf{v}_{i})-f_{\text{MLP}_{1}}(\mathbf{u}_{j})\Big\|$
$\leq\sqrt{d}\text{Lip}(f_{\text{MLP}_{2}})\text{Lip}(f_{\text{MLP}_{1}})\max_{i\in[m]}\min_{j\in[n]}\|\mathbf{v}_{i}-\mathbf{u}_{j}\|$
@@ B.7.3 The $\textsc{NN}_{\textsc{max}}$ model is Lipschitz continuous with respect to the Hausdorff distance
B.8 Proof of Lemma 3.5
B.8.1 The $\textsc{NN}_{\textsc{mean}}$ model is Lipschitz continuous with respect to the matching distance
Let $\mathcal{X}$ denote a set that contains multisets of vectors of equal cardinalities, i. e., $|X|=M$ and $\in\mathcal{S}(\mathbb{R}^{d}), X𝒳\forall$ $\in\mathcal{X}$ where $M\in\mathbb{N}$. Let $X,Y\in\mathcal{X}$ denote two multisets. We have shown in subsection B.7.1 above that:
$\bigg\|f_{\text{MLP}_{2}}\bigg(\frac{1}{m}\sum_{i=1}^{m}f_{\text{MLP}_{1}}(\mathbf{v}_{i})\bigg)-f_{\text{MLP}_{2}}\bigg(\frac{1}{n}\sum_{j=1}^{n}f_{\text{MLP}_{1}}(\mathbf{u}_{j})\bigg)\bigg\|$ $\leq\text{Lip}(f_{\text{MLP}_{1}})\text{Lip}(f_{\text{MLP}_{2}})d_{\text{EMD}}(X,Y)$
@@ B.8.1 The $\textsc{NN}_{\textsc{mean}}$ model is Lipschitz continuous with respect to the matching distance
$\bigg\|f_{\text{MLP}_{2}}\bigg(\frac{1}{m}\sum_{i=1}^{m}f_{\text{MLP}_{1}}(\mathbf{v}_{i})\bigg)-f_{\text{MLP}_{2}}\bigg(\frac{1}{n}\sum_{j=1}^{n}f_{\text{MLP}_{1}}(\mathbf{u}_{j})\bigg)\bigg\|$ $\leq\text{Lip}(f_{\text{MLP}_{1}})\text{Lip}(f_{\text{MLP}_{2}})d_{\text{EMD}}(X,Y)$
@@
B.8.2 The $\textsc{NN}_{\textsc{sum}}$ model is Lipschitz continuous with respect to the matching distance
Let $\mathcal{X}$ denote a set that contains multisets of vectors of equal cardinalities, i. e., $|X|=M$ and $\in\mathcal{S}(\mathbb{R}^{d}), X𝒳\forall$ $\in\mathcal{X}$ where $M\in\mathbb{N}$. Let $X,Y\in\mathcal{X}$ denote two multisets. Let $\pi^{*}$ denote the matching produced by the solution of the matching distance function:
$\pi^{*}=\operatorname*{arg\,min}_{\pi\in\mathfrak{S}_{M}}\sum_{i=1}^{M}\|\mathbf{v}_{\pi(i)}-\mathbf{u}_{i}\|\\$
@@ B.8.2 The $\textsc{NN}_{\textsc{sum}}$ model is Lipschitz continuous with respect to the matching distance
$\bigg\|f_{\text{MLP}_{2}}\bigg(\sum_{i=1}^{M}f_{\text{MLP}_{1}}(\mathbf{v}_{i})\bigg)-f_{\text{MLP}_{2}}\bigg(\sum_{j=1}^{M}f_{\text{MLP}_{1}}(\mathbf{u}_{j})\bigg)\bigg\|$ $\leq\text{Lip}(f_{\text{MLP}_{2}})\bigg\|\sum_{i=1}^{M}f_{\text{MLP}_{1}}(\mathbf{v}_{i})-\sum_{j=1}^{M}f_{\text{MLP}_{1}}(\mathbf{u}_{j})\bigg\|$
$=\text{Lip}(f_{\text{MLP}_{2}})\Big\|\sum_{i=1}^{M}\big(f_{\text{MLP}_{1}}(\mathbf{v}_{\pi^{*}(i)})-f_{\text{MLP}_{1}}(\mathbf{u}_{i})\big)\Big\|$
$\leq\text{Lip}(f_{\text{MLP}_{2}})\sum_{i=1}^{M}\Big\|f_{\text{MLP}_{1}}(\mathbf{v}_{\pi^{*}(i)})-f_{\text{MLP}_{1}}(\mathbf{u}_{i})\Big\|$
$\leq\text{Lip}(f_{\text{MLP}_{2}})\sum_{i=1}^{M}\text{Lip}(f_{\text{MLP}_{1}})\|\mathbf{v}_{\pi^{*}(i)}-\mathbf{u}_{i}\|$
$\leq\text{Lip}(f_{\text{MLP}_{1}})\text{Lip}(f_{\text{MLP}_{2}})\sum_{i=1}^{M}\|\mathbf{v}_{\pi^{*}(i)}-\mathbf{u}_{i}\|$
@@
which concludes the proof. The $\textsc{NN}_{\textsc{sum}}$ model is thus Lipschitz continuous with respect to the matching distance when the inputs are restricted to multisets of a fixed size, and its Lipschitz constant is upper bounded by $\text{Lip}(f_{\text{MLP}_{1}})\text{Lip}(f_{\text{MLP}_{2}})$.
B.8.3 The $\textsc{NN}_{\textsc{sum}}$ model is Lipschitz continuous with respect to EMD
Let $\mathcal{X}$ denote a set that contains multisets of vectors of equal cardinalities, i. e., $|X|=M$ and $\in\mathcal{S}(\mathbb{R}^{d}), X𝒳\forall$ $\in\mathcal{X}$ where $M\in\mathbb{N}$. Let $X,Y\in\mathcal{X}$ denote two multisets. We have shown in subsection B.8.2 above that:
$\bigg\|f_{\text{MLP}_{2}}\bigg(\sum_{i=1}^{M}f_{\text{MLP}_{1}}(\mathbf{v}_{i})\bigg)-f_{\text{MLP}_{2}}\bigg(\sum_{j=1}^{M}f_{\text{MLP}_{1}}(\mathbf{u}_{j})\bigg)\bigg\|$ $\leq\text{Lip}(f_{\text{MLP}_{1}})\text{Lip}(f_{\text{MLP}_{2}})d_{M}(X,Y)$